Ruthenium-106 (Ru106) is used to treat melanoma in the eye. This isotope decays by β−emission with a half-life of 373.59 days. One source of the isotope is reprocessed nuclear reactor fuel.

How many protons and how many neutrons does the Ru106 nucleus contain?

protons:

neutrons:

Identify the missing daughter nucleus in the β− emission decay of Ru106.

Ru106⟶

????

+ e−+ν¯e+ e−+ν¯e

How many years after Ru106 is implanted in the eye does it take for its decay rate (or activity) to be reduced to 26.5% of the initial decay rate? Use 1 year=365 days.

time:

year
Answer #1

(i) How many protons and neutrons does the 106Ru nucleus contain?

We know that, Number of proton is equal to the number of atomic number.

Nprotons = Z

where, Z = atomic number of 106Ru = 44

then, we get

Nprotons = 44

We know that, Nprotons + Nneutrons = A

Nneutrons = A - Nprotons

where, A = mass number of 106Ru = 106

then, we get

Nneutrons = [(106) - (44)]

Nneutrons = 62

(ii) Identify the missing daughter nucleus in the - emission decay of 106Ru.

106Ru44aXb + e- + e

To balance an above equation & we get

44 = (b - 1) b = 45

a = 106

Then, the missing daughter nucleus in the - emission decay of 106Ru will be given as :

106Rh45

Where, Rh = rhodium

(iii) How many years after 106Ru is implanted in the eye?

using a formula, we have

N = N0 e-t

we know that, = decay constant = ln (2) / 1/2

N = N0 e-(0.693)t/1/2

[(0.265) N0] / N0 = e-(0.693)t/1/2

ln (0.265) = - (0.693) t / 1/2

- (1.32803) = - (0.693) t / 1/2

where, 1/2 = half-life of isotope 106Ru = 373.59 days

then, we get

(0.693) t = [(1.32803) (373.59 days)]

t = [(496.1 days) / (0.693)]

t = 715.8 days

converting days into year :

t = (715.8 days) [(1 year) / (365 days)]

t = 1.96 year